Three heater coils H1, H2 and H3 rated (1000 W :
100V), (1500 W : 100 V) and (2000 W : 100 V) individually take a time interval of 12 minutes, 6 minutes and 4 minutes respectively to generate the same quantity of heat Q when connected to a 100 V source. Now these three coils are connected in parallel and this parallel combination is connected to a 50 V source. The time interval taken by this parallel combination of coils to generate the same quantity of heat Q is
  • 2 minutes
  • 6 minutes
  • 8 minutes
  • 12 minutes
  •  

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
H = I2RT  or H = P*t 
H1 = 1000* 12/60 = 200 joules 
H2 = 1500 * 6/60 = 150 joules 
H3 = 2000 * 4/60 = 400/3 joules 
Heat generated will be different for each coil ,